OpenStudy (anonymous):

Given the linear equation y= -7/8x+2, find the y- coordinates of the points (-8,),(0,) and (8,). Show all work please. Plot points and graph the linear equation.

OpenStudy (anonymous):

just plug in the values of x: e.g. y = -7/8 * -8 + 2 = 7 + 2 = 9 note i have assumes -7/8 = 'minus 7/8 time x'